• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Meziani

    Log entry time 08:01:03 on December 5,2007

    Entry number 218417

    keyword=Shift Summary


    Shift Begins
    00:00
    Change momentum setting on both arms
    p0=642.76 MeV/c dp/p = 0%
    Target 12 C (Carbon Elastic Calibration)
    00:13-00:18 Run # 3340 evts: 1.9M Collimators Open.

    Move target to Optics target

    10uA, dp/p=0%
    Runs # 3341 (left,SSlit OUT) evts:2.5M
    21481 (right,SSlit OUT) evts:2.8M

    50uA Sieve Slits In dp/p=0%
    Runs # 3342 (left,SSlit IN) evts: 1.58M
    21482(right,SSlit IN) evts: 1.8M

    Changing momentum on both arms to
    p0=629.91MeV/c dp/p = -2%
    50uA
    01:10-01:16 Run # 3343 (left, SSlit IN) evts:1.5M
    10uA
    01:24-01:24 Run # 3344 deadtime too high
    01:26-01:34 Run # 3345 (left SSLit OUT) evts:2.4M
    01:30-01:47 Run # 21483 (right,SSLit IN)evts:1.5M
    01:40-01:46 Run # 3346 (left,SSlit,OUT) evts:1.9M
    01:58-02:07 Run # 21484 (right,SSlit,OUT)evts:2.7M
    02:04-02:07 Run # 3347 (left, SSlit,IN) evts:0.3M
    50uA
    02:16-02:25 Run # 3348 (left,SSlit OUT)

    The rest of runs are summarized in separate elogs.
    Finished Optics runs and calibration runs (see elog)
    Finished scan of beam on empty target

    06:10 Both spectrometers at set at 40 degrees.
    15uA
    Start the scan with the 208 Pb target.
    Check the reconstructed target. Changed the momentum of the left arm
    to p0=0.61705 GeV/c.
    Scanning the lead target

    07:25 Changing the momentum to 330. MeV and taking data for
    the crystal calibration.

    Shift Ends
    <\pre>